# Global Bifurcation of Non-Radial Solutions for Symmetric Sub-linear Elliptic Systems on the Planar Unit Disc

### v1

@@ 4.1. Computation of the Local Bifurcation Invariant
Under the assumptions ($A_{1}$)($A_{5}$), conditions ($B_{1}$) and ($B_{2}$) are satisfied for the operator equation (8) (cf. Remarks 1.1, 4.1). Therefore, the existence of a branch of non-trivial solutions to (1) bifurcating from an isolated critical point $(\alpha_{0},0)\in\Lambda$ is reduced, by Theorem 3.1, to computation of the local bifurcation invariant $\omega_{G}(\alpha_{0})$. In addition, we assume ($A_{0}$), so that the local bifurcation invariant is well-defined at each critical point in $\Lambda$. Adopting notations from Section 3, choose $\alpha^{\pm}_{0}\in(\alpha_{0}-\epsilon,\alpha_{0}+\epsilon)$ with $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$, where $\epsilon>0$ is chosen such that, for all $0<|\alpha-\alpha_{0}|<\epsilon$, the solution $(\alpha,0)\in M$ is a regular point and put $\mathscr{A}(\alpha^{\pm}_{0}):=D_{u}\mathscr{F}(\alpha^{\pm}_{0},0)$. Then, the local bifurcation invariant $\omega_{G}(\alpha_{0})$ at the isolated critical point $(\alpha_{0},0)\in\mathbb{R}\times\mathscr{H}$ is given by
@@
where $B(\mathscr{H}):=\{u\in\mathscr{H}\;:\;\|u\|_{\mathscr{H}}<1\}$ is the open unit ball in $\mathscr{H}$. Notice that, since $(\alpha^{\pm}_{0},0)\in\mathbb{R}\times\mathscr{H}$ are regular points of (8), computation of the local bifurcation invariant $\omega_{G}(\alpha_{0})$ amounts to computation of the Leray-Schuader $G$-equivariant degree of the $G$-equivariant linear isomorphism,
@@
Under the assumptions ($A_{1}$)-($A_{5}$) and ($A_{0}$) and using the notation (19), the local bifurcation invariant at an isolated critical point $(\alpha_{0},0)\in\Lambda$ with deleted regular neighborhood $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$ on which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism is given by
@@
In order to formulate the main local equivariant bifurcation result, we must first consider some additional properties of the basic degree. For a more thorough exposition of these topics we refer the reader to Appendix C. Take $s\in\mathbb{N}$ and define the $s$-folding map as the Lie group homomorphism,
@@
An orbit type which is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ is also maximal in $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$. Therefore, any $u\in\mathscr{H}\setminus\{0\}$ with an isotropy $G_{u}\leq G$ such that $(G_{u})$ is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be radially symmetric. It follows that in order to detect branches of solutions to (7) corresponding to maps which are both non-trivial and non-radial, we must restrict our focus to orbit types which are maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m\in\mathbb{N}$. Indeed, to demonstrate the existence of a branch of non-radial solutions bifurcating from some isolated critical point $(\alpha_{0},0)\in\Lambda$, it is sufficient to show that for some $m>0$ there is an orbit type $(H)\in\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ with $\operatorname{coeff}^{H}(\omega_{G}(\alpha_{0}))\not=0$. Moreover, in such a case one can conclude that for some $s\geq 1$ there exists a branch non-radial solutions to (7) with symmetries at least $(^{s}H)$ bifurcating from $(\alpha_{0},0)$.
For $m>0$, denote by $\mathfrak{M}_{m}$ the set of all maximal orbit types in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ and by $\mathfrak{M}_{m,j}$ the set of orbit types $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})\cap\mathfrak{M}_{m}$. Since each $(H)\in\mathfrak{M}_{m}$ is also an orbit type in $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for at least one $j\in\{1,2,\ldots,r\}$, one has
@@
Notice that $\Psi_{s}(\mathfrak{M}_{m,j})=\mathfrak{M}_{sm,j}$ for any $m>0,\;s\in\mathbb{N}$ and $j\in\{1,2,\ldots,r\}$ (in particular, one has $\Psi_{s}(\mathfrak{M}_{1,j})=\mathfrak{M}_{s,j}$). Hence, any orbit type $(H_{0})\in\mathfrak{M}_{m,j}$ can be recovered from an orbit type in $\mathfrak{M}_{1,j}$ by the relation $(H):=\Psi_{s}^{-1}(H_{0})$.
@@
We need one more result, this concerning the fate of orbit types belonging to $\mathfrak{M}_{m}$ in the Burnside Ring product of basic degrees such as (20). In particular, we find that for $m>0$ and $i,l\in\{1,2,\ldots,r\}$, the coefficient of $(H)\in\mathfrak{M}_{m,i}\cap\mathfrak{M}_{m,l}$ is $2$-nilpotent with respect to the Burnside Ring product $\deg_{\mathcal{V}_{m,i}}\cdot\deg_{\mathcal{V}_{m,l}}\in A(G)$ in the case that both $\dim\mathcal{V}_{m,i}^{H}$ and $\dim\mathcal{V}_{m,l}^{H}$ are odd.
@@
$=(G)-(y_{i}+y_{l}-y_{i}y_{l}|W(H)|)(H)+a,$
where $a\in A(G)$ is such that $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\{1,2,\ldots\}$, and put $y_{0}:=y_{i}+y_{l}-y_{i}y_{l}|W(H)|$. Now, if $\dim\mathcal{V}_{1,i}^{H}$ and $\dim\mathcal{V}_{1,l}^{H}$ are both even, then one has $y_{i}=y_{l}=0$ and the result follows. On the other hand, if $\dim\mathcal{V}_{1,i}^{H}$ and $\dim\mathcal{V}_{1,l}^{H}$ are both odd, then one has $y_{i}=y_{l}=x_{0}$ such that
@@
Let $(\alpha_{0},0)\in\Lambda$ be an isolated critical point with a deleted regular neighborhood $\alpha_{0}^{-}<\alpha_{0}<\alpha_{0}^{+}$ on which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism and for a given $s\in\mathbb{N}$, $(H)\in\mathfrak{M}_{1}$ put
@@
Take $(H)\in\mathfrak{M}_{1}$, $(n,m,j)\in\Sigma$ and $s\in{\mathbb{N}}$. If $\operatorname{coeff}^{{}^{s}H}(\deg_{\mathcal{V}_{m,j}})\neq 0$, then $m=s$. On the other hand, if $({}^{s}H)\leq({}^{m}H)$, then $s\mid m$.
In particular, we are interested in keeping track of the numbers $s\in{\mathbb{N}}$ for which the parities of $\mathfrak{n}^{s}(\alpha_{0}^{\pm},H)$ disagree. With this in mind, put
@@
If $(\alpha_{0},0)\in\Lambda$ is an isolated critical point with a deleted regular neighborhood $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$ on which the linear operator $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism and if there is an orbit type $(H)\in\mathfrak{M}_{1}$ and a number $\mathfrak{s}:=\mathfrak{s}(\alpha_{0},H)\in{\mathbb{N}}$ such that $\mathfrak{i}^{\mathfrak{s}}(\alpha_{0},H)\not=0$ and $\mathfrak{i}^{m}(\alpha_{0},H)=0$ for all $m>\mathfrak{s}$, then
@@
For convenience, take $\alpha^{*}_{0}\in\{\alpha^{\pm}_{0}\}$ and introduce the notation
@@
such that the local bifurcation invariant (cf. 3.1) becomes
@@
and, for each $s=0,1,\ldots$ define
(31) $\rho^{s}_{G}(\alpha^{*}_{0},H):=\prod\limits_{(n,m,j)\in\Sigma^{s}(\alpha^{*}_{0},H)}\deg_{\mathcal{V}_{m,j}}.$
Clearly the index set (19) has the partition
$\Sigma(\alpha_{0}^{*})=\bigcup\limits_{s\in\mathbb{N}\cup\{0\}}\Sigma^{s}(\alpha_{0}^{*},H),$
such that the product of basic degrees (30) becomes
$\rho_{G}(\alpha^{*}_{0})=\prod\limits_{s\in\mathbb{N}\cup\{0\}}\rho^{s}_{G}(\alpha^{*}_{0},H).$
It follows, from Lemma (4.3) and its Corollary (4.1), that each of (31) is of the form
@@
Now, since $\mathfrak{n}^{s}(\alpha^{\pm}_{0},H)$ are of the same parity for all $s>\mathfrak{s}$, one has
$\rho_{G}^{s}(\alpha^{-}_{0})\cdot\rho_{G}^{s}(\alpha^{+}_{0})=(G)+c_{s},$
@@
(32) =±$=\pm x_{0}({}^{s_{0}}H)\mp y_{s}(\alpha^{*})x_{0}({}^{s}H)\cdot({}^{s_{0}}H)+\bm{\alpha}$
where $\bm{\alpha}\in A(G)$ is such that $\operatorname{coeff}^{{}^{s^{\prime}}H}(\bm{\alpha})=0$ for all $s^{\prime}\in\mathbb{N}$. Now, if $\mathfrak{n}^{s}(\alpha^{*}_{0},H)$ is even, then $y_{s}(\alpha^{*}_{0})=0$ and the result follows. Suppose instead that $2\nmid\mathfrak{n}^{s}(\alpha^{*}_{0},H)$. Then (4.1) becomes
@@ 4.2. Resolution of the Rabinowitz Alternative
Without an appropriate fixed point reduction of the bifurcation problem (8), one is unable to guarantee that a branch of non-trivial solutions $\mathcal{C}$ to (1) bifurcating from a given isolated critical point $(\alpha_{0},0)\in\mathbb{R}\times\mathscr{H}$ whose existence has been established using a Krasnosel’skii type result (e.g. by Theorem 4.1) is not comprised of radial solutions. With this in mind, consider the subgroup $\bm{K}:=\{(-1,e,-1),(1,e,1)\}\leq O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ and denote by $\mathscr{F}^{\bm{K}}:\mathbb{R}\times\mathscr{H}^{\bm{K}}\to\mathscr{H}^{\bm{K}}$ the restriction
@@
is also solution to (8). Notice also that any radial solution to (34) belongs to the set of trivial solutions
$M^{\bm{K}}:=\{(\alpha,0):\alpha\in\mathbb{R}\times\mathscr{H}^{\bm{K}}\}.$
Therefore, any branch of of non-trivial solutions to the bifurcation problem (34) consists solely of non-radial solutions. In this $\bm{K}$-fixed point setting, we adapt each of the notions introduced in Section 3.2 used to describe the Rabinowitz alternative for the equation (8) to the bifurcation map (33). To begin, notice that the $\bm{K}$ fixed point space $\mathscr{H}^{\bm{K}}$ is an isometric Hilbert representation of the group $\mathcal{G}:=N(\bm{K})/\bm{K}=O(2)\times\Gamma$ with the $\mathcal{G}$-isotypic decomposition
@@
We denote by $\mathscr{S}^{\bm{K}}$ the set of non-trivial solutions to (34), i.e.
@@
As before, the critical set of (34), now denoted $\Lambda^{\bm{K}}$, is the set of trivial solutions $(\alpha_{0},0)\in M^{\bm{K}}$ for which $\mathscr{A}^{\bm{K}}(\alpha_{0})$ is not an isomorphism
$\Lambda^{\bm{K}}:=\{(\alpha,0)\in\mathbb{R}\times\mathscr{H}^{\bm{K}}:\mathscr{A}^{\bm{K}}(\alpha):\mathscr{H}^{\bm{K}}\to\mathscr{H}^{\bm{K}}\;\;\text{ is not an isomorphism}\}.$
@@
(Rabinowitz’ Alternative) Under the assumptions ($A_{0}$)($A_{5}$), let $\mathcal{U}\subset\mathbb{R}\times\mathscr{H}^{\bm{K}}$ be an open bounded $\mathcal{G}$-invariant set with $\partial\mathcal{U}\cap\Lambda^{\bm{K}}=\emptyset$ and $(\alpha_{0},0)\in\mathcal{U}\cap\Lambda^{\bm{K}}$. If $\mathcal{C}$ is a branch of nontrivial solutions to (34) bifurcating from the critical point $(\alpha_{0},0)$, then one has the following alternative:

### v2

@@ 4.1. Computation of the Local Bifurcation Invariant
Under the assumptions ($A_{1}$)($A_{5}$), conditions ($B_{1}$) and ($B_{2}$) are satisfied for the operator equation (8) (cf. Remarks 1.1, 4.1). Therefore, the existence of a branch of non-trivial solutions to (1) bifurcating from an isolated critical point $(\alpha_{0},0)\in\Lambda$ is reduced, by Theorem 3.1, to computation of the local bifurcation invariant $\omega_{G}(\alpha_{0})$. Adopting the notations from Section 3, choose $\alpha^{\pm}_{0}\in(\alpha_{0}-\epsilon,\alpha_{0}+\epsilon)$ with $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$, where $\epsilon>0$ is chosen such that, for all $0<|\alpha-\alpha_{0}|<\epsilon$, the solution $(\alpha,0)\in M$ is a regular point and put $\mathscr{A}(\alpha^{\pm}_{0}):=D_{u}\mathscr{F}(\alpha^{\pm}_{0},0)$. Then, the local bifurcation invariant $\omega_{G}(\alpha_{0})$ at the isolated critical point $(\alpha_{0},0)\in\mathbb{R}\times\mathscr{H}$ is given by
@@
where $B(\mathscr{H}):=\{u\in\mathscr{H}\;:\;\|u\|_{\mathscr{H}}<1\}$ is the open unit ball in $\mathscr{H}$. Notice that, since $(\alpha^{\pm}_{0},0)\in\mathbb{R}\times\mathscr{H}$ are regular points of (8), computation of the local bifurcation invariant $\omega_{G}(\alpha_{0})$ amounts to computation of the Leray-Schuader $G$-equivariant degree of the $G$-equivariant linear isomorphism
@@
Under the assumptions ($A_{0}$)-($A_{5}$) and using the notation (20), the local bifurcation invariant at an isolated critical point $(\alpha_{0},0)\in\Lambda$ with deleted regular neighborhood $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$ on which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism is given by
@@
In order to formulate the main local equivariant bifurcation result, we must first consider some additional properties of the basic degree. For a more thorough exposition of these topics we refer the reader to [4]. Take $s\in\mathbb{N}$ and define the $s$-folding map as the Lie group homomorphism,
@@
An orbit type which is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ is also maximal in $\Phi_{0}(G;\mathscr{H}_{0}\setminus\{0\})$. Therefore, any $u\in\mathscr{H}\setminus\{0\}$ with an isotropy $G_{u}\leq G$ such that $(G_{u})$ is maximal in $\Phi_{0}(G;\mathscr{H}\setminus\{0\})$ must be radially symmetric. It follows that in order to detect branches of solutions to (7) corresponding to maps which are both non-trivial and non-radial, we must restrict our focus to orbit types which are maximal in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ for some positive $m\in\mathbb{N}$. Indeed, to demonstrate the existence of a branch of non-radial solutions bifurcating from some isolated critical point $(\alpha_{0},0)\in\Lambda$, it is sufficient to show that for some $m>0$ there is an orbit type $(H)\in\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ with $\operatorname{coeff}^{H}(\omega_{G}(\alpha_{0}))\not=0$.In such a case, one can additionally conclude that for some $s\geq 1$ there exists a branch non-radial solutions to (7) with symmetries at least $(^{s}H)$ bifurcating from $(\alpha_{0},0)$.
With motivation from Remark 4.2, we denote by $\mathfrak{M}_{m}$, $m>0$ the set of all maximal orbit types in $\Phi_{0}(G;\mathscr{H}_{m}\setminus\{0\})$ and by $\mathfrak{M}_{m,j}$ the set of orbit types $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})\cap\mathfrak{M}_{m}$. Since each $(H)\in\mathfrak{M}_{m}$ is also an orbit type in $\Phi_{0}(G;\mathscr{H}_{m,j}\setminus\{0\})$ for at least one $j\in\{1,2,\ldots,r\}$, one has
@@
Notice that $\Psi_{s}(\mathfrak{M}_{m,j})=\mathfrak{M}_{sm,j}$ for any $s\in\mathbb{N}$ and $j\in\{1,2,\ldots,r\}$ (in particular, one has $\Psi_{s}(\mathfrak{M}_{1,j})=\mathfrak{M}_{s,j}$). Hence, any orbit type $(H_{0})\in\mathfrak{M}_{m,j}$ can be recovered from an orbit type in $\mathfrak{M}_{1,j}$ by the relation $(H):=\Psi_{s}^{-1}(H_{0})$.
@@
The following result concerns the fate of orbit types belonging to $\mathfrak{M}_{m}$ in the Burnside Ring product of basic degrees such as (21). In particular, we find that for $m>0$ and $i,l\in\{1,2,\ldots,r\}$, the coefficient of $(H)\in\mathfrak{M}_{m,i}\cap\mathfrak{M}_{m,l}$ is $2$-nilpotent with respect to the Burnside Ring product $\deg_{\mathcal{V}_{m,i}}\cdot\deg_{\mathcal{V}_{m,l}}\in A(G)$ in the case that both $\dim\mathcal{V}_{m,i}^{H}$ and $\dim\mathcal{V}_{m,l}^{H}$ are odd.
@@
$=(G)-(y_{i}+y_{l}-y_{i}y_{l}|W(H)|)({}^{m}H)+a,$
where $a\in A(G)$ is such that $\operatorname{coeff}^{{}^{s}H}(a)=0$ for all $s\in\{1,2,\ldots\}$, and put $y_{0}:=y_{i}+y_{l}-y_{i}y_{l}|W(H)|$, i.e.
$y_{0}=\operatorname{coeff}^{{}^{m}H}(\deg_{\mathcal{V}_{m,i}}\cdot\deg_{\mathcal{V}_{m,l}}).$
Now, if $\dim\mathcal{V}_{1,i}^{H}$ and $\dim\mathcal{V}_{1,l}^{H}$ are both even, then one has $y_{i}=y_{l}=0$ and the result follows. On the other hand, if $\dim\mathcal{V}_{1,i}^{H}$ and $\dim\mathcal{V}_{1,l}^{H}$ are both odd, then one has $y_{i}=y_{l}=x_{0}$ such that
@@
Let $(\alpha_{0},0)\in\Lambda$ be an isolated critical point with a deleted regular neighborhood $\alpha_{0}^{-}<\alpha_{0}<\alpha_{0}^{+}$ on which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism. For a given $s\in\mathbb{N}$, $(H)\in\mathfrak{M}_{1}$ we put
@@
Take $(H)\in\mathfrak{M}_{1}$, $(n,m,j)\in\Sigma$ and $s\in{\mathbb{N}}$. If $\operatorname{coeff}^{{}^{s}H}(\deg_{\mathcal{V}_{m,j}})\neq 0$, then $m=s$. On the other hand, if $({}^{s}H)\leq({}^{m}H)$, then $s\mid m$. The converse of each statement is not, in general, true.
In order to keep track of the numbers $s\in{\mathbb{N}}$ for which the parities of $\mathfrak{n}^{s}(\alpha_{0}^{\pm},H)$ disagree, we put
@@
If $(\alpha_{0},0)\in\Lambda$ is an isolated critical point with a deleted regular neighborhood $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$ on which the linear operator $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism and if there is an orbit type $(H)\in\mathfrak{M}_{1}$ such that $\mathfrak{i}^{\mathfrak{s}}(\alpha_{0},H)\not=0$ and $\mathfrak{i}^{m}(\alpha_{0},H)=0$ for all $m>\mathfrak{s}$, where $\mathfrak{s}:=\mathfrak{s}(\alpha_{0},H)\in{\mathbb{N}}$, then one has
@@
For convenience, take $\alpha^{*}_{0}\in\{\alpha^{\pm}_{0}\}$. Adopting the notation
@@
the local bifurcation invariant (cf. 3.1) becomes
@@
and, for each $s=0,1,\ldots$ defines
(32) $\rho^{s}_{G}(\alpha^{*}_{0},H):=\prod\limits_{(n,m,j)\in\Sigma^{s}(\alpha^{*}_{0},H)}\deg_{\mathcal{V}_{m,j}},$
then the product of basic degrees (31) becomes
$\rho_{G}(\alpha^{*}_{0})=\prod\limits_{s\in\mathbb{N}\cup\{0\}}\rho^{s}_{G}(\alpha^{*}_{0},H),$
since $\Sigma(\alpha_{0}^{*})=\bigcup_{s\in\mathbb{N}\cup\{0\}}\Sigma^{s}(\alpha_{0}^{*},H)$ is a partition of the index set (20). It follows, from Lemma (4.3) and its Corollary (4.1), that each of (32) is of the form
@@
Since $\mathfrak{n}^{s}(\alpha^{\pm}_{0},H)$ are of the same parity for all $s>\mathfrak{s}$, one has
$\rho_{G}^{s}(\alpha^{-}_{0})\cdot\rho_{G}^{s}(\alpha^{+}_{0})=(G)+c_{s},\;s>\mathfrak{s},$
@@
(33) =±$=\pm x_{0}({}^{s_{0}}H)\mp y_{s}(\alpha^{*}_{0})x_{0}({}^{s}H)\cdot({}^{s_{0}}H)+\bm{\alpha}$
where $\bm{\alpha}\in A(G)$ is such that $\operatorname{coeff}^{{}^{s^{\prime}}H}(\bm{\alpha})=0$ for all $s^{\prime}\in\mathbb{N}$. Now, if $\mathfrak{n}^{s}(\alpha^{*}_{0},H)$ is even, then $y_{s}(\alpha^{*}_{0})=0$ and the result follows. Supposing instead that $2\nmid\mathfrak{n}^{s}(\alpha^{*}_{0},H)$, then (4.1) becomes
@@ 4.2. Resolution of the Rabinowitz Alternative
Without an appropriate fixed point reduction of the bifurcation problem (8), we are unable to guarantee that a branch of non-trivial solutions $\mathcal{C}$ to (1) bifurcating from a given isolated critical point $(\alpha_{0},0)\in\mathbb{R}\times\mathscr{H}$ whose existence has been established using a Krasnosel’skii type result (e.g. by Theorem 4.1) is not comprised of radial solutions. With this in mind, consider the subgroup $\bm{K}:=\{(-1,e,-1),(1,e,1)\}\leq O(2)\times\Gamma\times{\mathbb{Z}}_{2}$ and denote by $\mathscr{F}^{\bm{K}}:\mathbb{R}\times\mathscr{H}^{\bm{K}}\to\mathscr{H}^{\bm{K}}$ the restriction
@@
is also solution to (8). Notice also that any radial solution to (35) belongs to the set of trivial solutions (cf. Condition ($B_{1}$))
[…]
Therefore, any branch of of non-trivial solutions to the bifurcation problem (35) consists solely of non-radial solutions. In this $\bm{K}$-fixed point setting, we must adapt each of the notions introduced in Section 3.2 used to describe the Rabinowitz alternative for the equation (8) to the bifurcation map (34). To begin, notice that the $\bm{K}$-fixed point space $\mathscr{H}^{\bm{K}}$ is an isometric Hilbert representation of the group $\mathcal{G}:=N(\bm{K})/\bm{K}=O(2)\times\Gamma$ with the $\mathcal{G}$-isotypic decomposition
@@
We denote by $\mathscr{S}^{\bm{K}}$ the set of $\bm{K}$-fixed non-trivial solutions to (8), i.e.
@@
As before, the critical set of (35), now denoted $\Lambda^{\bm{K}}$, is the set of trivial solutions $(\alpha_{0},0)\in M$ for which $\mathscr{A}^{\bm{K}}(\alpha_{0})$ is not an isomorphism
$\Lambda^{\bm{K}}:=\{(\alpha,0)\in\mathbb{R}\times\mathscr{H}:\mathscr{A}^{\bm{K}}(\alpha):\mathscr{H}^{\bm{K}}\to\mathscr{H}^{\bm{K}}\;\;\text{ is not an isomorphism}\}.$
@@
(Rabinowitz’ Alternative) Under the assumptions ($A_{0}$)($A_{5}$), let $\mathcal{U}\subset\mathbb{R}\times\mathscr{H}^{\bm{K}}$ be an open bounded $\mathcal{G}$-invariant set with $\partial\mathcal{U}\cap\Lambda^{\bm{K}}=\emptyset$. If $\mathcal{C}$ is a branch of nontrivial solutions to (35) bifurcating from an isolated critical critical point $(\alpha_{0},0)\in\mathcal{U}\cap\Lambda$, then one has the following alternative: